Three boxes of equal dimensions (20 cm x 30cm x 50 cm), containing material which density 6.8x10-2 g/cm3 are to be stored for winter. What is the least amount of pressure that can be exerted by the blocks?
a) 1999.2 g/cm.s2
b) 40 Pa
c) 666 Pa
d) 4000 g/cm.s2
e) 333 Pa
1
Expert's answer
2019-07-22T11:09:48-0400

Mass of the one block:

"m = \rho V"

Volume of one block:

"V = S*h"

According to the condition:

"h = 50cm = 0.5m"

Pressure force:

"F = P = mg"

Respectively:

"F = P = \rho Vg = \rho Shg"

Determine the pressure:

"p = F/S = {\rho Shg \over S} = \rho gh"

We obtain the initial formula for calculating the pressure:

"p = 68(kg\/m^3)*9.8(N\/kg)*0.5(m) = 333.2(N\/m^2) = 333.2(Pa)"

"333.2Pa \approx 333Pa"

Answer: e) 333 Pa